# Self-Data Distillation for Recovering Quality in Pruned Large Language Models

### v1

@@ Abstract
Large language models have driven significant progress in natural language processing, but their deployment requires substantial compute and memory resources. As models scale, compression techniques become essential for balancing model quality with computational efficiency. Structured pruning, which removes less critical components of the model, is a promising strategy for reducing complexity. However, one-shot pruning often results in significant quality degradation, particularly in tasks requiring multi-step reasoning. To recover lost quality, supervised fine-tuning (SFT) is commonly applied, but it can lead to catastrophic forgetting by shifting the model’s learned data distribution. Therefore, addressing the degradation from both pruning and SFT is essential to preserve the original model’s quality. In this work, we propose self-data distilled fine-tuning to address these challenges. Our approach leverages the original, unpruned model to generate a distilled dataset that preserves semantic richness and mitigates catastrophic forgetting by maintaining alignment with the base model’s knowledge. Empirically, we demonstrate that self-data distillation consistently outperforms standard SFT, improving average accuracy by up to 8% on the HuggingFace OpenLLM Leaderboard v1. Specifically, when pruning 6 decoder blocks on Llama3.1-8B Instruct (i.e., 32 to 24 layers, reducing the model size from 8.03B to 6.72B parameters), our method retains 91.2% of the original model’s accuracy compared to 81.7% with SFT, while reducing real-world FLOPs by 16.30%. Furthermore, our approach scales effectively across datasets, with the quality improving as the dataset size increases.
@@ 2.1 Layer-Pruning Algorithm for Language Models
Figure 2: Comparison of Llama3.1-8B Instruct using (left) angular cosine and (center) block influence (BI) score metrics. (right) Distribution of embedding similarities on GSM8k after fine-tuning on OpenMathInstruct for a pruned Llama3.1-8B Instruct model (i.e., 24 decoder layers, prune block size $n$ = 6). Self-data distilled fine-tuning (Self-Data FT) achieves higher similarity to the original baseline, indicating reduced distribution shift compared to standard supervised fine-tuning (SFT).
@@ Appendix C Experimental Setup for Understanding Catastrophic Forgetting
To understand the impact of distribution shift on catastrophic forgetting, we conducted experiments using the baseline model (i.e., Llama3.1-8B Instruct) and its pruned variants fine-tuned with both supervised fine-tuning (SFT) and self-data distilled fine-tuning (Self-Data FT). Specifically, we pruned 6 decoder layers, reducing the model from 32 to 24 layers, and evaluated the models on the GSM8k dataset. For these experiments, we generated model responses using the baseline and pruned variants on the GSM8k dataset to capture how the distribution shift affects reasoning tasks post-pruning. Following Yang et al. (2024), to quantify the distribution shift, we employed Sentence-BERT (Reimers and Gurevych, 2019) to derive sentence embeddings from the model-generated responses. Then, similar to the method proposed by Zhang et al. (2023), we calculated the cosine similarity between the sentence embeddings of the pruned models and those generated by the original Llama3.1-8B Instruct model.
@@ D.1 Geometric Consistency and Application
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
[…]
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
[…]
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
[…]
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
[…]
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [No]
[…]
  21. Guidelines:
  22. 6.
    Experimental Setting/Details
  23. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  24. Answer: [Yes]
[…]
  26. 7.
    Experiment Statistical Significance
  27. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  28. Answer: [Yes]
[…]
  30. 8.
    Experiments Compute Resources
  31. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  32. Answer: [Yes]
[…]
  34. 9.
    Code Of Ethics
  35. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  36. Answer: [Yes]
[…]
  38. 10.
    Broader Impacts
  39. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  40. Answer: [No]
[…]
  42. 11.
    Safeguards
  43. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  44. Answer: [N/A]
[…]
  46. 12.
    Licenses for existing assets
  47. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  48. Answer: [Yes]
[…]
  50. 13.
    New Assets
  51. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  52. Answer: [N/A]
[…]
  54. 14.
    Crowdsourcing and Research with Human Subjects
  55. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  56. Answer: [N/A]
[…]
  58. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  59. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  60. Answer: [N/A]
[…]

### v2

@@ Abstract
Large language models have driven significant progress in natural language processing, but their deployment requires substantial compute and memory resources. As models scale, compression techniques become essential for balancing model quality with computational efficiency. Structured pruning, which removes less critical components of the model, is a promising strategy for reducing complexity. However, one-shot pruning often results in significant quality degradation, particularly in tasks requiring multi-step reasoning. To recover lost quality, supervised fine-tuning (SFT) is commonly applied, but it can lead to catastrophic forgetting by shifting the model’s learned data distribution. Therefore, addressing the degradation from both pruning and SFT is essential to preserve the original model’s quality. In this work, we propose self-data distilled fine-tuning to address these challenges. Our approach leverages the original, unpruned model to generate a distilled dataset that preserves semantic richness and mitigates catastrophic forgetting by maintaining alignment with the base model’s knowledge. Empirically, we demonstrate that self-data distillation consistently outperforms standard SFT, improving average accuracy by up to 8% on the HuggingFace OpenLLM Leaderboard v1. Specifically, when pruning 6 decoder blocks on Llama3.1-8B Instruct (i.e., 32 to 26 layers, reducing the model size from 8.03B to 6.72B parameters), our method retains 91.2% of the original model’s accuracy compared to 81.7% with SFT, while reducing real-world FLOPs by 16.30%. Furthermore, our approach scales effectively across datasets, with the quality improving as the dataset size increases.
@@ 2.1 Layer-Pruning Algorithm for Language Models
Figure 2: Comparison of Llama3.1-8B Instruct using (left) angular cosine and (center) block influence (BI) score metrics. (right) Distribution of embedding similarities on GSM8k after fine-tuning on OpenMathInstruct for a pruned Llama3.1-8B Instruct model (i.e., 26 decoder layers, prune block size $n$ = 6). Self-data distilled fine-tuning (Self-Data FT) achieves higher similarity to the original baseline, indicating reduced distribution shift compared to standard supervised fine-tuning (SFT).
@@ Appendix C Experimental Setup for Understanding Catastrophic Forgetting
To understand the impact of distribution shift on catastrophic forgetting, we conducted experiments using the baseline model (i.e., Llama3.1-8B Instruct) and its pruned variants fine-tuned with both supervised fine-tuning (SFT) and self-data distilled fine-tuning (Self-Data FT). Specifically, we pruned 6 decoder layers, reducing the model from 32 to 26 layers, and evaluated the models on the GSM8k dataset. For these experiments, we generated model responses using the baseline and pruned variants on the GSM8k dataset to capture how the distribution shift affects reasoning tasks post-pruning. Following Yang et al. (2024), to quantify the distribution shift, we employed Sentence-BERT (Reimers and Gurevych, 2019) to derive sentence embeddings from the model-generated responses. Then, similar to the method proposed by Zhang et al. (2023), we calculated the cosine similarity between the sentence embeddings of the pruned models and those generated by the original Llama3.1-8B Instruct model.